you read a report stating that correctional officers have very low job satisfaction. you take 3 samples of n=4 correctional officers who work in either a supermax,medium. and minimum security prison and ask them to self-report their level of stress on a 1-20point scale. the data are presented below Supermax: 17, 14, 13, 12, T1=56 SS1=14 meidum: 12, 14, 11, 13, T2=50 SS2=5 Minimum: 10, 8, 13, 9, T3=40 SS3=14 G=146 EX2=1842
Answer #1

This a one factor-ANOVA testing

Supermax Medium Minimum
17 12 10
14 14 8
13 11 13
12 13 9

Ho: All are equal VS H1 All u are not equal

F - Test Statistic = \frac{MSS_{B}}{MSS_{R}}

SSB MSSB 1 = \frac{32.667}{2}=16.33    (Variance between the treatments/ No. of treatment - 1)

MSS_{R}=\frac{SS_{R}}{n-k}\frac{33}{9}=3.667 (Variance between the residual / Total obs. - no. of treatment)

SS_{B}=\sum \frac{y_{i.}^{2}}{n_{i}}-\frac{G^{2}}{N} = 32.667 (y_{i}. is the total of individual treatment scores eg. sum of supermax and G is the total of all the scores)

SS_{R}=SS_{T}-SS_{B} = 33

SS_{T}=\sum \sum y_{ij}^{2}-\frac{G^{2}}{N} = 65.667 (\sum \sum y_{ij}^{2} is the sum of individual squares of  scores)

Variance df SS MS
Between 2 32.67 16.33
Residual 9 33 3.667
Total 11 65.667

F - Stat = \frac{16.33}{3.667} =4.455

Using 5% level of significance the critical value

F_{\alpha\%/2,k-1,n-k}=F_{2.5\%,2,9}=4.2565

F - Stat > Critical Value

Decision : We reject the null hypothesis and conclude that there is a significant different between the level of stress under different level of correctional officers.